# Irreducibility of polynomials with random multiplicative coefficients revisited

### v1

@@ Acknowledgments
The second author is supported by the project “Group schemes, root systems, and related representations” founded by the European Union - NextGenerationEU through Romania’s National Recovery and Resilience Plan (PNRR) call no. PNRR-III-C9-2023- I8, Project CF159/31.07.2023, and coordinated by the Ministry of Research, Innovation and Digitalization (MCID) of Romania. The authors would like to thank Lior Bary-Soroker and David Hokken for the comments on a preliminary version of this manuscript. The authors would like to thank the Oberwolfach Mathematics Institute for hosting them and the organizers of the workshops (K. Matomäki, J. Maynard , K. Soundararajan, T.D. Wooley and L. Bary-Soroker, A. Ostafe, and P. Sarnak) that facilitated this collaboration.
@@ 2. Proof of Main Theorem
Consider $\tilde{P}_{f,N}(X)=P_{f,N}(X+1)$ where $N=2^{k}$ and $j\geq 1$ and fix large $A\ll\log\log N.$ Then

### v2

@@ 1. Introduction
In the last section we briefly discuss application of this result to the deterministic family of Turyn polynomials.
@@ Acknowledgments
The second author is supported by the project “Group schemes, root systems, and related representations” founded by the European Union - NextGenerationEU through Romania’s National Recovery and Resilience Plan (PNRR) call no. PNRR-III-C9-2023- I8, Project CF159/31.07.2023, and coordinated by the Ministry of Research, Innovation and Digitalization (MCID) of Romania. THe authors would like to thank Lior Bary-Soroker and David Hoken for the comments on a preliminary version of this manuscript. The authors would like to thank the Oberwolfach Mathematics Institute for hosting them and the organizers of the workshops (K. Matomäki, J. Maynard , K. Soundararajan, T.D. Wooley and L. Bary-Soroker, A. Ostafe, and P. Sarnak) that fascilitated this collaboration.
@@ 2. Proof of Main Theorem
Consider $\tilde{P}_{f,N}(X)=P_{f,N}(X+1)=\sum_{j\leq N}a_{j}x^{j}$ where $N=2^{k}$ and $j\geq 1$ and fix large $A\ll\log\log N.$ Then
@@
2.1. Irreducibility of Turyn polynomials
As mentioned in the introduction, one motivating example for considering polynomials with random multiplicative coefficients is the study of the irreducibility of the Turyn polynomials given by
$F_{d,p}(z)=\sum_{a=1}^{d}\left(\frac{a}{p}\right)z^{a},$
see Problem 1 in [36]. These polynomials are of significant interest in number theory and analysis, and they play an important role in extremal problems for Littlewood polynomials; see [9], [23], [32] for some results and for extensive lists of references. Varjú and Xu [36] proved that, under the assumption of ERH, if a prime $p$ is drawn uniformly at random from the set $[d,h(d)]$ with $h(d)\geq 2^{\pi(d)}d^{4}$, then
$\lim_{p\rightarrow\infty}\mathbb{P}[F_{d,p}(X)\hskip 5.69054pt\text{is irreducible}]=1.$
Since the distribution of the sequence $\left(\frac{a}{p}\right)$ for $a\leq d$ and $p\geq C^{d}$ is well modeled by random multiplicative functions, one can directly apply our Theorem 1.1 together with unconditional variants of the character sum bounds from [36] to obtain the following unconditional corollary.
Corollary 2.4.
Let $d=2^{k}.$ Then, if a prime $p$ is drawn uniformly at random from the set $[d,h(d)]$ with $h(d)\geq 2^{2^{\pi(d)}}$, we have
$\lim_{k\rightarrow\infty}\mathbb{P}[F_{d,p}(X)\hskip 5.69054pt\text{is irreducible}]=1.$